# Self-embedding similitudes of Bedford-McMullen carpets with dependent ratios

### v2

@@ 4. The dependent self-affine case
Let $K,m,n$ be as at the beginning of this section. If $f\in\mathcal{S}(\mathbb{R}^{2})$ sends $K$ into itself, then $f$ is not oblique.
Proof.
By Remark 3.3 and Proposition 2.2, it suffices to consider when $N\leq n-1$ and #$\#J\leq m-1$. Suppose on the contrary that $f$ is oblique. Pick $g,C$ as in Lemma 4.1 with $g$ oblique and denote by $\rho$ the contraction ratio of $g$. If $N=1$, then it is not hard to see that every horizontal slice of $K$ is a finite set. From the form of $C$ (recall (4.5)), $C$ is a finite set. But by Proposition 2.1 and the obliqueness of $g$, $\pi_{1}(g(K))$ is an infinite set. This contradicts $\pi_{1}(g(K))\subset\pi_{1}(C\times\mathbb{R})=C$.
Now suppose $N\geq 2$. By Remark 4.2 (see (4.6)), we can find $z\in\pi_{2}(K)$, $c\in\mathbb{R}$, $k\geq 1$ and a level-$k$ rectangle $R$ such that
(4.7) $g\varphi_{R}(K^{y}\times\{y\})\subset(K^{z}+c)\times\mathbb{R},\quad\forall y\in\pi_{2}(K).$
@@
(4.8) $g\varphi_{R}(K^{y}\times\{y\})\subset L_{y}\cap((K^{z}+c)\times\mathbb{R}),\quad\forall y\in\pi_{2}(K).$
@@
Note that for each $y\in\pi_{2}(K)$, $g\varphi_{R}(K^{y}\times\{y\})$ can be regarded as a similar copy of $K^{y}$ contracted by $n^{-k}\rho$, while $L_{y}\cap((K^{z}+c)\times\mathbb{R})$ can be regarded as a translated copy of $\frac{|v|}{v_{1}}K^{z}$. In particular, one can obtain from (4.8) a similitude $h_{y}\in\mathcal{S}(\mathbb{R})$ with contraction ratio $n^{-k}\rho\cdot\frac{v_{1}}{|v|}$ (which is independent of $y$) such that
@@
Pick $j_{*}\in J$ so that #$\#I_{j_{*}}=N$ and let $y_{*}:=\sum_{k=1}^{\infty}j_{*}m^{-k}$. Then $K^{y_{*}}=E(n,I_{j_{*}})$ (recall Section 2.2). By (4.9), $h_{y_{*}}(K^{y_{*}})\subset K^{z}$. So $0<\mathcal{H}^{\alpha}(K^{z})<\infty$, where $\alpha:=\frac{\log N}{\log n}$. It might be suitable for us to point out an observation here (which will be of help later but not in this proof): since (4.9) also implies $h_{z}(K^{z})\subset K^{z}$, it follows from Proposition 2.6 that
@@
Next, since $h_{z}(K^{z})\subset K^{z}$, by Corollary 2.7, $h_{z}$ maps a left isolated point $a$ and a right one $b$ both to one-sided isolated points of $K^{z}$. Since $h_{z}(K^{z})\subset K^{z}$ is simply a one-dimensional reformulation of the inclusion $g\varphi_{R}(K^{z}\times\{z\})\subset L_{z}\cap((K^{z}+c)\times\mathbb{R})$, this in turn gives us two points $\widetilde{a},\widetilde{b}\in g\varphi_{R}(K^{z}\times\{z\})$ and some $\delta>0$ such that both of the tubes $(\pi_{1}(\widetilde{a})-\delta,\pi_{1}(\widetilde{a}))\times\mathbb{R}$ and $(\pi_{1}(\widetilde{b}),\pi_{1}(\widetilde{b})+\delta)\times\mathbb{R}$ do not intersect $(K^{z}+c)\times\mathbb{R}$. Moreover, there are two level-$t$ rectangles $R_{t},R^{\prime}_{t}\subset R$ such that $\widetilde{a}\in g\varphi_{R_{t}}(K)$ and $\widetilde{b}\in g\varphi_{R^{\prime}_{t}}(K)$, respectively, and (again by Corollary 2.7) $\pi_{2}(R_{t})=\pi_{2}(R^{\prime}_{t})$ for all $t\geq k_{0}$ for some $k_{0}$. See Figure 5 for an illustration.
@@
Pick $x,x^{\prime}\in K$ with $\pi_{2}(x)=\max\pi_{2}(K)$ and $\pi_{2}(x^{\prime})=\min\pi_{2}(K)$. For convenience, let $e_{1}:=(1,0)$ and $e_{2}:=(0,1)$. Since $\widetilde{a}\in g\varphi_{R_{t}}(K)$, $\widetilde{b}\in g\varphi_{R^{\prime}_{t}}(K)$, $x$ can be expressed as
(4.11) $x=(g\varphi_{R_{t}})^{-1}(\widetilde{a})+c_{t}e_{1}+d_{t}e_{2}=(g\varphi_{R^{\prime}_{t}})^{-1}(\widetilde{b})+\gamma_{t}e_{1}+\eta_{t}e_{2}$
@@
(4.12) $x^{\prime}=(g\varphi_{R_{t}})^{-1}(\widetilde{a})+c^{\prime}_{t}e_{1}+d^{\prime}_{t}e_{2}=(g\varphi_{R^{\prime}_{t}})^{-1}(\widetilde{b})+\gamma^{\prime}_{t}e_{1}+d^{\prime}_{t}e_{2}.$
@@
So at least one of $d_{t},-d^{\prime}_{t}$ is no less than $\beta/2$. Writing $e_{1,1}:=\pi_{1}(g(e_{1}))$ and $e_{2,1}:=\pi_{1}(g(e_{2}))$, we have by the obliqueness of $g$ that $e_{1,1},e_{2,1}\neq 0$. We distinguish four simple cases.
Case 1: $d_{t}\geq\beta/2$ and $e_{2,1}>0$
@@
$g\varphi_{R^{\prime}_{t}}(x)$ $=\widetilde{b}+g\varphi_{R^{\prime}_{t}}(\gamma_{t}e_{1})+g\varphi_{R^{\prime}_{t}}(d_{t}e_{2})$
$=\widetilde{b}+n^{-t}\gamma_{t}g(e_{1})+m^{-t}d_{t}g(e_{2}).$
Since $d_{t}\geq\beta/2$, $n>m$ and $|\gamma_{t}|\leq 1$, $0<n^{-t}\gamma_{t}e_{1,1}+m^{-t}d_{t}e_{2,1}<\delta$ for all large $t$ and hence
$\pi_{1}(g\varphi_{R^{\prime}_{t}}(x))=\pi_{1}(\widetilde{b})+n^{-t}\gamma_{t}e_{1,1}+m^{-t}d_{t}e_{2,1}\in(\pi_{1}(\widetilde{b}),\pi_{1}(\widetilde{b})+\delta).$
@@
$g\varphi_{R_{t}}(x)$ $=\widetilde{a}+g\varphi_{R_{t}}(c_{t}e_{1})+g\varphi_{R_{t}}(d_{t}e_{2})$
$=\widetilde{a}+n^{-t}c_{t}g(e_{1})+m^{-t}d_{t}g(e_{2}).$
Since $d_{t}\geq\beta/2$, $n>m$ and $|c_{t}|\leq 1$, we have $-\delta<n^{-t}c_{t}e_{1,1}+m^{-t}d_{t}e_{2,1}<0$ for all large $t$ and hence
$\pi_{1}(g\varphi_{R_{t}}(x))=\pi_{1}(\widetilde{a})+n^{-t}c_{t}e_{1,1}+m^{-t}d_{t}e_{2,1}\in(\pi_{1}(\widetilde{a})-\delta,\pi_{1}(\widetilde{a})).$
@@
Case 3: $d^{\prime}_{t}\geq\beta/2$ and $e_{2,1}>0$
Applying the argument in Case 1 to (4.12), we have for all large $t$ that
$\pi_{1}(g\varphi_{R^{\prime}_{t}}(x^{\prime}))=\pi_{1}(\widetilde{b})+n^{-t}\gamma^{\prime}_{t}e_{1,1}+m^{-t}d^{\prime}_{t}e_{2,1}\in(\pi_{1}(\widetilde{b}),\pi_{1}(\widetilde{b})+\delta),$
which leads to a contradiction.
Case 4: $d^{\prime}_{t}\geq\beta/2$ and $e_{2,1}<0$
Applying the argument in Case 2 to (4.12), we have for all large $t$ that
$\pi_{1}(g\varphi_{R_{t}}(x^{\prime}))=\pi_{1}(\widetilde{a})+n^{-t}c^{\prime}_{t}e_{1,1}+m^{-t}d^{\prime}_{t}e_{2,1}\in(\pi_{1}(\widetilde{a})-\delta,\pi_{1}(\widetilde{a})),$
which leads to a contradiction. ∎
Combining with the above result and Proposition 2.6, we can quickly prove Theorem 1.5.
@@
Let $K,n,m$ be as at the beginning of this section. If $f\in\mathcal{S}(\mathbb{R}^{2})$ sends $K$ into itself, then the above theorem tells us that $f$ is not oblique. So it is not hard to find some integer $k\geq 1$ such that $f^{k}(x)=\lambda^{k}x+\eta$ for some $\eta\in\mathbb{R}^{2}$, i.e., the orthogonal part of $f^{k}$ is simply the identity matrix.
Note that $f^{k}$ sends the $x$-axis to a horizontal line. If $\frac{\log\lambda}{\log n}\notin\mathbb{Q}$, we can apply Lemma 4.1 to $f^{k}$ and get (recall the second paragraph in Remark 4.2) some $g\in\mathcal{S}(\mathbb{R}^{2})$ of which the contraction ratio $\rho$ is an irrational power of $n$. However, if $N\geq 2$, applying the argument in the proof of the above theorem (where now $v=(1,0)$) , we can still arrive at (4.10) and hence $\frac{\log\rho}{\log n}\in\mathbb{Q}$. This is a contradiction.
If $N=1$, then every row contains at most one selected rectangle in the initial pattern of $K$. In particular, #$\#\Lambda=\#J\leq m<n$. So there must be a vacant column. On the other hand, since #$\#\Lambda\geq 2$ and $K$ is not contained in any vertical line, #$\#I\geq 2$. Therefore, $\pi_{1}(K)=E(n,I)$ is a deleted-digit set with Hausdorff dimension $10<\frac{\log\#I}{\log n}\leq\frac{\log\#\Lambda}{\log n}<$ 1. Noting that $f^{k}(K)\subset K$ implies that
$\lambda^{k}(E(n,I))+\pi_{1}(\eta)=\lambda^{k}(\pi_{1}(K))+\pi_{1}(\eta)=\pi_{1}(f^{k}(K))\subset\pi_{1}(K)=E(n,I).$
Then it follows from Lemma 2.5 that $\frac{\log\lambda}{\log n}\in\mathbb{Q}$. This completes the proof. ∎

### v3

@@ 4. The dependent self-affine case
If $f$ is a similitude sending $K$ into itself, then $f$ is not oblique.
Proof of Theorem 4.3: the contracting case.
By Proposition 2.2, it suffices to consider when $N\leq n-1$. Suppose on the contrary that $f$ is oblique. Pick $g,C$ as in Lemma 4.1 with $g$ oblique and denote by $\rho$ the contraction ratio of $g$. If $N=1$, then it is not hard to see that every horizontal slice of $K$ is a finite set. From the form of $C$ (recall (4.5)), $C$ is a finite set. But by Proposition 2.1 and the obliqueness of $g$, $\pi_{1}(g(K))$ is an infinite set. This contradicts $\pi_{1}(g(K))\subset\pi_{1}(C\times\mathbb{R})=C$.
Now suppose $N\geq 2$. By Remark 4.2 (see (4.6)), we can find $z\in\pi_{2}(K)$, $i\in\mathbb{Z}$, $k\geq 1$ and a level-$k$ rectangle $R$ such that
(4.7) $g\varphi_{R}(K^{y}\times\{y\})\subset(K^{z}+i)\times\mathbb{R},\quad\forall y\in\pi_{2}(K).$
@@
(4.8) $g\varphi_{R}(K^{y}\times\{y\})\subset L_{y}\cap((K^{z}+i)\times\mathbb{R}),\quad\forall y\in\pi_{2}(K).$
@@
Note that for each $y\in\pi_{2}(K)$, $g\varphi_{R}(K^{y}\times\{y\})$ can be regarded as a similar copy of $K^{y}$ contracted by $n^{-k}\rho$, while $L_{y}\cap((K^{z}+i)\times\mathbb{R})$ can be regarded as a translated copy of $\frac{|v|}{v_{1}}K^{z}$. In particular, one can obtain from (4.8) a similitude $h_{y}\in\mathcal{S}(\mathbb{R})$ with contraction ratio $n^{-k}\rho\cdot\frac{v_{1}}{|v|}$ (which is independent of $y$) such that
@@
Pick $j_{*}\in J$ so that #$\#I_{j_{*}}=N$ and let $y_{*}:=\sum_{k=1}^{\infty}j_{*}m^{-k}$. Then $K^{y_{*}}=E(n,I_{j_{*}})$ (recall Section 2.2). By (4.9), $h_{y_{*}}(K^{y_{*}})\subset K^{z}$. So $0<\mathcal{H}^{\alpha}(K^{z})<\infty$, where $\alpha:=\frac{\log N}{\log n}$ (recall the observation before Lemma 2.3). It might be suitable for us to point out an observation here (which will be of help later but not in this proof): since (4.9) also implies $h_{z}(K^{z})\subset K^{z}$, it follows from Proposition 2.6 that
@@
Next, since $h_{z}(K^{z})\subset K^{z}$, by Corollary 2.7, $h_{z}$ maps a left isolated point $a$ and a right one $b$ both to one-sided isolated points of $K^{z}$. Since $h_{z}(K^{z})\subset K^{z}$ is simply a one-dimensional reformulation of the inclusion $g\varphi_{R}(K^{z}\times\{z\})\subset L_{z}\cap((K^{z}+i)\times\mathbb{R})$, this in turn gives us two points $\widetilde{a},\widetilde{b}\in g\varphi_{R}(K^{z}\times\{z\})$ and some $\delta>0$ such that both of the tubes $(\pi_{1}(\widetilde{a})-\delta,\pi_{1}(\widetilde{a}))\times\mathbb{R}$ and $(\pi_{1}(\widetilde{b}),\pi_{1}(\widetilde{b})+\delta)\times\mathbb{R}$ do not intersect $(K^{z}+i)\times\mathbb{R}$. Moreover, there are two level-$t$ rectangles $R_{t},R^{\prime}_{t}\subset R$ such that $\widetilde{a}\in g\varphi_{R_{t}}(K)$ and $\widetilde{b}\in g\varphi_{R^{\prime}_{t}}(K)$, respectively, and (again by Corollary 2.7) $\pi_{2}(R_{t})=\pi_{2}(R^{\prime}_{t})$ for all $t\geq k_{0}$ for some $k_{0}$. See Figure 5 for an illustration.
@@
Pick $x,x^{\prime}\in K$ with $\pi_{2}(x)=\max\pi_{2}(K)$ and $\pi_{2}(x^{\prime})=\min\pi_{2}(K)$. Since $\widetilde{a}\in g\varphi_{R_{t}}(K)$, $\widetilde{b}\in g\varphi_{R^{\prime}_{t}}(K)$, $x$ can be expressed as
(4.11) $x=(g\varphi_{R_{t}})^{-1}(\widetilde{a})+\Big{(}\begin{matrix}c_{t}\\ d_{t}\end{matrix}\Big{)}=(g\varphi_{R^{\prime}_{t}})^{-1}(\widetilde{b})+\Big{(}\begin{matrix}\gamma_{t}\\ \eta_{t}\end{matrix}\Big{)}$
@@
(4.12) $x^{\prime}=(g\varphi_{R_{t}})^{-1}(\widetilde{a})+\Big{(}\begin{matrix}c^{\prime}_{t}\\ d^{\prime}_{t}\end{matrix}\Big{)}=(g\varphi_{R^{\prime}_{t}})^{-1}(\widetilde{b})+\Big{(}\begin{matrix}\gamma^{\prime}_{t}\\ \eta^{\prime}_{t}\end{matrix}\Big{)}.$
@@
So at least one of $d_{t},-d^{\prime}_{t}$ is no less than $\beta/2$. Writing $O=(\begin{smallmatrix}o_{1,1}&o_{1,2}\\ o_{2,1}&o_{2,2}\end{smallmatrix})$ to be the orthogonal part of $g$, we have by the obliqueness of $g$ that $o_{1,1},o_{1,2}\neq 0$. We distinguish four simple cases.
Case 1: $d_{t}\geq\beta/2$ and $o_{1,2}>0$
@@
$g\varphi_{R^{\prime}_{t}}(x)-\widetilde{b}=\Big{(}\begin{matrix}o_{1,1}&o_{1,2}\\ o_{2,1}&o_{2,2}\end{matrix}\Big{)}\cdot\Big{(}\begin{matrix}\rho n^{-t}\gamma_{t}\\ \rho m^{-t}d_{t}\end{matrix}\Big{)}.$
Since $d_{t}\geq\beta/2$, $n>m$ and $|\gamma_{t}|\leq 1$, $0<\rho n^{-t}\gamma_{t}o_{1,1}+\rho m^{-t}d_{t}o_{1,2}<\delta$ when $t$ is large and hence
$\pi_{1}(g\varphi_{R^{\prime}_{t}}(x))=\pi_{1}(\widetilde{b})+\rho n^{-t}\gamma_{t}o_{1,1}+\rho m^{-t}d_{t}o_{1,2}\in(\pi_{1}(\widetilde{b}),\pi_{1}(\widetilde{b})+\delta).$
@@
$g\varphi_{R_{t}}(x)-\widetilde{a}=\Big{(}\begin{matrix}o_{1,1}&o_{1,2}\\ o_{2,1}&o_{2,2}\end{matrix}\Big{)}\cdot\Big{(}\begin{matrix}\rho n^{-t}c_{t}\\ \rho m^{-t}d_{t}\end{matrix}\Big{)}.$
Since $d_{t}\geq\beta/2$, $n>m$ and $|c_{t}|\leq 1$, we have $-\delta<\rho n^{-t}c_{t}o_{1,1}+\rho m^{-t}d_{t}o_{1,2}<0$ when $t$ is large and hence
$\pi_{1}(g\varphi_{R_{t}}(x))=\pi_{1}(\widetilde{a})+\rho n^{-t}c_{t}o_{1,1}+\rho m^{-t}d_{t}o_{1,2}\in(\pi_{1}(\widetilde{a})-\delta,\pi_{1}(\widetilde{a})).$
@@
Case 3: $-d^{\prime}_{t}\geq\beta/2$ and $o_{1,2}>0$
Applying the argument in Case 2 to (4.12), we have
$\pi_{1}(g\varphi_{R_{t}}(x^{\prime}))=\pi_{1}(\widetilde{a})+\rho n^{-t}c^{\prime}_{t}o_{1,1}+\rho m^{-t}d^{\prime}_{t}o_{1,2}\in(\pi_{1}(\widetilde{a})-\delta,\pi_{1}(\widetilde{a}))$
when $t$ is large, which leads to a contradiction.
Case 4: $-d^{\prime}_{t}\geq\beta/2$ and $o_{1,2}<0$
Applying the argument in Case 1 to (4.12), we have
$\pi_{1}(g\varphi_{R^{\prime}_{t}}(x^{\prime}))=\pi_{1}(\widetilde{b})+\rho n^{-t}\gamma^{\prime}_{t}o_{1,1}+\rho m^{-t}d^{\prime}_{t}o_{1,2}\in(\pi_{1}(\widetilde{b}),\pi_{1}(\widetilde{b})+\delta)$
when $t$ is large, which leads to a contradiction. ∎
For the isometry case, we need a rescaling lemma of which the argument is analogous to the one of Lemma 4.1.
Lemma 4.4.
Let $f\in\mathcal{S}(\mathbb{R}^{2})$ be an oblique isometry. If $f(K)\subset K$, then we can find some $z\in\pi_{2}(K)$ and a nondegenerate scaled copy $E$ of $E(m,J)$ such that $E\subset K^{z}$.
Proof.
Pick any sequence $\{R_{k}\}_{k=1}^{\infty}$, where each $R_{k}$ is a level-$k$ rectangle and $R_{k+1}\subset R_{k}$ for all $k\geq 1$. Since $f$ is an isometry, we have
$\max\{|\pi_{1}(f(R_{k}))|,|\pi_{2}(f(R_{k}))|\}<2m^{-k},\quad\forall k\geq 1.$
As a result, $f(R_{pk})$ is contained in at most $4$ squares of side length $m^{-pk}=n^{-qk}$. More precisely, there are $x\in n^{-qk}\mathbb{Z},y\in m^{-pk}\mathbb{Z}$ such that
$f\varphi_{R_{pk}}(K)\subset(x,x+2n^{-qk})\times(y,y+2m^{-pk})=:Q_{k}.$
Denote by $h_{k}$ the homothety sending $(0,1)^{2}$ to the square $(x,x+n^{-qk})\times(y,y+m^{-pk})=:Q^{\prime}_{k}$, the “left bottom square” of $Q_{k}$.
Since $E(m,J)=\pi_{2}(K)$, it is easy to see that $d_{H}(\varphi_{R_{pk}}(K),\varphi_{R_{p_{k}}}(\{0\}\times E(m,J)))\leq n^{-pk}$, where $d_{H}$ denotes the Hausdorff distance (see [8]). So
(4.13) $d_{H}(h_{k}^{-1}f\varphi_{R_{pk}}(K),h_{k}^{-1}f\varphi_{R_{pk}}(\{0\}\times E(m,J)))\leq n^{(q-p)k}\to 0,\quad k\to\infty.$
Moreover, writing $O$ to be the orthogonal part of $f$, each $h_{k}^{-1}f\varphi_{R_{pk}}(\{0\}\times E(m,J))$ is simply a translated copy of $O(\{0\}\times E(m,J))$, say $O(\{0\}\times E(m,J))+a_{k}$. So (4.13) becomes
$d_{H}(h_{k}^{-1}f\varphi_{R_{pk}}(K),O(\{0\}\times E(m,J))+a_{k})\leq n^{(q-p)k}\to 0,\quad k\to\infty.$
By the definition of $h_{k}$, the sequence $\{a_{k}\}_{k}$ lies in a bounded region in the plane. For notational simplicity, let us assume without loss of generality that $\{a_{k}\}_{k}$ is convergent, say $a_{k}\to a$. Then under the Hausdorff distance,
$h_{k}^{-1}f\varphi_{R_{pk}}(K)\to O(\{0\}\times E(m,J))+a,\quad k\to\infty.$
Applying the rescaling argument in the proof of Lemma 4.1 (now to $\{h_{k}^{-1}f\varphi_{R_{pk}}\}$ instead of $\{h_{t}^{-1}f^{t}\}_{t}$), we can find at most four points $z_{i,j}$, $0\leq i,j\leq 1$, such that
$O(\{0\}\times E(m,J))+a\subset\bigcup_{0\leq i,j\leq 1}(K^{z_{i,j}}+i)\times\mathbb{R}.$
Since $O$ is oblique, projecting the above two sets to the $x$-axis, we see that $\bigcup_{0\leq i,j\leq 1}(K^{z_{i,j}}+i)$ contains a nondegenerate scaled copy of $E(m,J)$. Then by Baire’s theorem, some $K^{z_{i,j}}+i$ must contain an interior part of that copy of $E(m,J)$. Since $E(m,J)$ is self-similar, this completes the proof. ∎
Proof of Theorem 4.3: the isometry case.
By Remark 3.3 and Proposition 2.2, it suffices to consider when $N\leq n-1$ and #$\#J\leq m-1$. Since $K$ is not contained in any line, #$\#J>1$. Thus $E(m,J)$ is a deleted-digit set with Hausdorff dimension $\frac{\log\#J}{\log m}\in(0,1)$.
Assume on the contrary that $f$ is oblique. Note that
$f(K^{y}\times\{y\})\subset f(K)\subset K\subset\mathbb{R}\times\pi_{2}(K)=\mathbb{R}\times E(m,J),\quad\forall y\in\pi_{2}(K).$
Therefore, $\pi_{2}(f(K^{y}\times\{y\}))\subset E(m,J)$ for all $y$. Since $f$ is oblique, the left hand side is a nondegenerate scaled copy of $K^{y}$. Picking $y_{*}$ as in the proof of Theorem 4.3 (before (4.10)), we have $\operatorname{dim_{H}}E(m,J)\geq\operatorname{dim_{H}}K^{y_{*}}=\frac{\log N}{\log n}=:\alpha$.
Pick $z\in\pi_{2}(K)$ according to the above lemma. So there are also $\beta_{2},c_{2}\in\mathbb{R}$ such that $E(m,J)\subset\beta_{2}K^{z}+c_{2}$. On the other hand, we have seen that $E(m,J)$ contains a scaled copy of $K^{z}$, say $\beta_{1}K^{z}+c_{1}$. So $\operatorname{dim_{H}}E(m,J)\leq\operatorname{dim_{H}}K^{z}\leq\alpha$ and hence $\operatorname{dim_{H}}E(m,J)=\alpha$. Since $\alpha>0$, $N\geq 2$. Since $\mathcal{H}^{\alpha}(E(m,J))>0$, $\mathcal{H}^{\alpha}(K^{z})\in(0,\infty)$ (again, the <absent<$\infty$< ∞ part is an observation before Lemma 2.3).
From $f(K^{z}\times\{z\})\subset\mathbb{R}\times E(m,J)\subset\mathbb{R}\times(\beta_{2}K^{z}+c_{2})$ and the above observation, we can apply Corollary 2.7 and argue as in the second part of the proof of Theorem 4.3 (the one after (4.10)) and obtain a contradiction (the only difference is that in the previous proof, we embedded some oblique slice into a union of vertical lines, whereas now we embed some oblique slice into a union of horizontal lines).∎
Combining with these results and Proposition 2.6, we can quickly prove Theorem 1.5.
@@
Let $K,n,m,f$ be as in Theorem 1.5. If $\lambda=1$ then there is nothing to prove. If $\lambda<1$, then Theorem 4.3 tells us that $f$ is not oblique. So it is easy to find some integer $k\geq 1$ such that $f^{k}(x)=\lambda^{k}x+\eta$ for some $\eta\in\mathbb{R}^{2}$, i.e., the orthogonal part of $f^{k}$ is simply the identity matrix.
Note that $f^{k}$ sends the $x$-axis to a horizontal line. If $\frac{\log\lambda}{\log n}\notin\mathbb{Q}$, we can apply Lemma 4.1 to $f^{k}$ and get (recall the second paragraph in Remark 4.2) some $g\in\mathcal{S}(\mathbb{R}^{2})$ of which the contraction ratio $\rho$ is an irrational power of $n$. Also, since $f^{k}$ involves no rotations nor reflections, so does $g$. However, if $12\leq N\leq n-$ 1, applying the argument in the proof of the above theorem (where now $v=(1,0)$), we can still arrive at (4.10) and hence $\frac{\log\rho}{\log n}\in\mathbb{Q}$. This is a contradiction.
If $N=1$, then every row contains at most one selected rectangle in the initial pattern of $K$. In particular, #$\#\Lambda=\#J\leq m<n$. So there must be a vacant column. On the other hand, since #$\#\Lambda\geq 2$ and $K$ is not contained in any vertical line, #$\#I\geq 2$. Therefore, $\pi_{1}(K)=E(n,I)$ is a deleted-digit set with Hausdorff dimension $10<\frac{\log\#I}{\log n}\leq\frac{\log\#\Lambda}{\log n}<$ 1. Since $f^{k}(K)\subset K$, $\pi_{1}(f^{k}(K))\subset\pi_{1}(K)=E(n,I)$. So
$\lambda^{k}E(n,I)+\pi_{1}(\eta)=\lambda^{k}\pi_{1}(K)+\pi_{1}(\eta)=\pi_{1}(f^{k}(K))\subset E(n,I).$
Then it follows from Lemma 2.5 that $\frac{\log\lambda}{\log n}\in\mathbb{Q}$.
Finally, assume that $N=n$. Let $J^{\prime}:=\{j\in J:\#I_{j}=n\}$. Note that
(4.14) $\{y\in\pi_{2}(K):K^{y}\text{ contains an interval}\}\subset\bigcup_{t\geq 0}\bigcup_{z\in\mathbb{Z}}\frac{E(m,J^{\prime})+z}{n^{t}}.$
The case when #$\#J^{\prime}=1$ has been settled in [3, Section 6.6]. If #$\#J^{\prime}\geq 2$ (clearly, #$\#J^{\prime}\leq m-1$), then $10<\operatorname{dim_{H}}E(m,J^{\prime})<$ 1. Note that $K^{y}$ is an interval of length $1$ whenever $y\in E(m,J^{\prime})$. Since $f^{k}(K)\subset K$, it follows from (4.14) that
$\lambda^{k}E(m,J^{\prime})+\pi_{2}(\eta)\subset\bigcup_{t\geq 0}\bigcup_{z\in\mathbb{Z}}\frac{E(m,J^{\prime})+z}{n^{t}}.$
By Baire’s theorem, there are some $t,z$ such that $\frac{E(m,J^{\prime})+z}{n^{t}}$ contains an interior part of $\lambda^{k}E(m,J^{\prime})+\pi_{2}(\eta)$. By Lemma 2.5, $\frac{\log\lambda}{\log n}\in\mathbb{Q}$, which completes the proof. ∎